# Real-Time Operator Takeover for Visuomotor Diffusion Policy Training

### v1

@@ Abstract
We present a Real-Time Operator Takeover (RTOT) paradigm enabling operators to seamlessly take control of a live visuomotor diffusion policy, guiding the system back into desirable states or reinforcing specific demonstrations. We presents new insights in using the Mahalonobis distance to automaicaly identify undesirable states. Once the operator has intervened and redirected the system, the control is seamlessly returned to the policy, which resumes generating actions until further intervention is required. We demonstrate that incorporating the targeted takeover demonstrations significantly improves policy performance compared to training solely with an equivalent number of, but longer, initial demonstrations. We provide an in-depth analysis of using the Mahalanobis distance to detect out-of-distribution states, illustrating its utility for identifying critical failure points during execution. Supporting materials, including videos of initial and takeover demonstrations and all rice-scooping experiments, are available on the project website111 Operator-Takeover.github.io.
@@ I Introduction
While teleoperation systems can boost the performance of IL methods, they do not address the fundamental problem of how to cover sufficient variations in demonstrations/tasks to make the learned policy more robust. Enhancing robustness in IL is an active area of research [15, 16]. Although the existing approaches improve the handling of scene variations, they still struggle when encountering states that are not present in the training data. Policies often fail in such scenarios, getting stuck due to the lack of demonstrations for recovering from undesirable states.
In this work, we tackle this challenge with our Real-Time Operator Takeover (RTOT) paradigm. The core idea is that even the most skilled operator may struggle to anticipate and demonstrate recovery scenarios or failure cases the policy might encounter. Our RTOT approach addresses this limitation by first training an initial policy with a smaller set of demonstrations than typically used in the literature. This policy is then deployed in the environment while the operator remains on standby. When the operator observes the system entering or approaching an undesirable state, they seamlessly take control to teleoperate the robot back to a desirable state. This takeover demonstrations are recorded, and a new policy is trained using both the initial and the newly recorded takeover demonstrations. As shown in Fig. 1, this paradigm can be repeated as many times as necessary to achieve the desired performance for any given task. Our contributions are:
@@
    Introduction and validation of the RTOT paradigm on a real-world rice-scooping task.
@@ IV Experimental Setup
We evaluate this through a cyclic rice scooping task, as shown in Fig. 1. The objective is to scoop rice from the bowl on the left (positioned in the white rectangle) and deposit it into the bowl on the right (which can be placed anywhere within its respective rectangle). To address the cyclic nature of the task, the initial demonstrations are designed to both start and end at approximately the same position after successfully completing one scoop.
To begin, we collect an initial dataset $\mathcal{D}_{I}^{60}$ consisting of 60 expert demonstrations. These demonstrations are performed while varying the locations of the rice bowls to introduce diversity in the training data. From this comprehensive dataset, we extract two smaller subsets: the first 20 demonstrations form $\mathcal{D}_{I}^{20}$, and the first 40 demonstrations form $\mathcal{D}_{I}^{40}$. Using these datasets, we train corresponding policies: $\pi_{I}^{20}$, $\pi_{I}^{40}$, and $\pi_{I}^{60}$.
Next, we deploy $\pi_{I}^{20}$ in the rice scooping environment as described in Section III. During this deployment, we collect 20 additional takeover demonstrations, denoted as $\mathcal{D}T^{20a}$, while systematically varying the rice bowl positions. These takeover demonstrations are then combined with the initial dataset $\mathcal{D}I^{20}$ to form an augmented dataset $\mathcal{D}{IT}^{a}$, which is used to train a new policy, $\pi{IT}^{a}$.
@@
Additionally, the combination of initial and takeover demonstrations helps to address failure cases more effectively, as takeover demonstrations specifically focus on challenging scenarios encountered during policy deployment. This iterative process not only improves policy robustness but also reduces the overall training time required to achieve high task performance.
@@ V Experimental Evaluation
We evaluate the performance of five different policies ($\pi_{I}^{20},\pi_{I}^{40},\pi_{IT}^{a},\pi_{I}^{60},\pi_{IT}^{b}$) on the cyclic rice scooping task. The goal of the task is to transfer as much rice as possible from a full bowl to an empty bowl within $45$ seconds, measured in grams. For each policy, we conduct $10$ trials, each with different placement locations for the bowls. Videos of all $50$ experiments are available on the project website.
@@
Results and Analysis. The results reveal a clear trend: the performance of baseline policies improves with the number of initial demonstrations. For example, $\pi_{I}^{20}$, trained on only $20$ demonstrations, achieved an average of $4.00$ grams of rice scooped across $10$ trials. Increasing the dataset to $40$ demonstrations boosted the performance significantly, with $\pi_{I}^{40}$ averaging $19.80$ grams. Using all $60$ initial demonstrations, $\pi_{I}^{60}$ achieved a substantial improvement, scooping an average of $41.00$ grams per trial. This highlights the importance of larger datasets for improving baseline performance in visuomotor policy training.
To address the question of whether employing the real-time operator takeover paradigm leads to better task performance, we compare the performance of the baseline $\pi_{I}^{40}$ with that of the takeover-enhanced policy $\pi_{IT}^{a}$. Remarkably, $\pi_{IT}^{a}$ achieved an average of $35.40$ grams scooped, representing a $79\%$ improvement over $\pi_{I}^{40}$. Similarly, the final policy, $\pi_{IT}^{b}$, trained on both initial and two sets of takeover demonstrations, outperformed $\pi_{I}^{60}$ by $20\%$, with an average of $49.20$ grams scooped per trial.
@@ VI To Takeover or Not to Takeover
In the current framework, the decision of when to perform a takeover is left entirely to the operator. We argue that this is a sensible approach, given that the quality of the initial demonstrations—and more broadly, the overall task performance—is highly influenced by the quality of these demonstrations [5]. However, having a more systematic way of assessing how close the robot is to undesirable states, or detecting when such states occur, could provide significant benefits.
@@
This motivates the need for a metric that evaluates how far observations obtained during inference deviate from the training distribution. Such a metric could serve as an out-of-distribution (OOD) measure to better inform intervention decisions.
@@ VI-A Mahalanobis Distance as an OOD Measure
We estimate the mean ($\mu$) and covariance matrix ($\Sigma$) of each embedding set using the Minimum Covariance Determinant Estimator implemented in scikit-learn [59]. For each encoded observation $h\in H$, the Mahalanobis distance $d_{M}$ is calculated using Algorithm 1.
@@ VI-B Mahalanobis Distance Analysis
However, an interesting anomaly is observed when comparing $\mathcal{D}I^{40}$ to $\mathcal{D}{IT}^{a}$ or $\mathcal{D}I^{60}$ to $\mathcal{D}{IT}^{b}$: the mean Mahalanobis distance does not directly correlate with task performance. For example, $\pi_{IT}^{b}$ outperforms $\pi_{I}^{60}$ in scooping rice but has a similar Mahalanobis distance profile.
Case Study: Trial 3. To explore this further, we analyze a specific trial in Fig. 7. For Trial number three, we compare the Mahalanobis distances of policies $\pi_{I}^{20}$, $\pi_{IT}^{b}$, and $\pi_{I}^{60}$. We observe that $\pi_{I}^{20}$ starts of with the highst distance while $\pi_{IT}^{b}$, and $\pi_{I}^{60}$ are both around $40$. the better in task performing $\pi_{IT}^{b}$ however then exhibits a slightly higher distance as the trail goes on however non of the states encounterd would need requiere the takeover of the operator and the distance stays below $100$ for the entiere duration of the trail.
The poorly performing policy $\pi_{I}^{20}$ exhibits the highest Mahalanobis distance peaking at around $160$, around the $25s$ mark, when the robot’s spoon moves outside the target bowl and fails to recover. The robot remains stuck in this state until the trial ends, highlighting the relationship between high Mahalanobis distance and undesirable states. In contrast, both $\pi_{I}^{60}$ and $\pi_{IT}^{b}$ maintain relatively lower Mahalanobis distances throughout the trial, with no significant increases. While $\pi_{IT}^{b}$ outperforms $\pi_{I}^{60}$ in task performance, the Mahalanobis distance does not differentiate much between the two policies in this regard. We also observe that the sudden stops and therefore camera shaking have significantly influence as can be observed on the last peak in $\pi_{IT}^{b}$ (orange), and $\pi_{I}^{60}$ (green) around the $45$s mark which was introduced by the sudden stop of the robot as the experiment was concluded. Furthermore, we see that the distance remains constant as long as no changes in the image are introduced via motion.
@@ VII Conclusion
In this work, we introduced a real-time operator takeover (RTOT) paradigm for training visuomotor diffusion policies. This framework empowers users to seamlessly take control of the robot from a suboptimal policy, preventing failures during execution. These takeover demonstrations are subsequently integrated into the training process, iteratively refining the policy until the desired performance is achieved.
Our experimental evaluation demonstrates the effectiveness of this approach. Despite the total time of takeover demonstrations being significantly shorter than that of the initial demonstrations, policies trained on these targeted datasets consistently outperform their counterparts. In a cyclic rice scooping task, policies enhanced with takeover demonstrations achieved superior task performance, highlighting the value of addressing failure cases during policy deployment rather than relying solely on extensive initial datasets.
Beyond showcasing the utility of the RTOT framework, we also explored the use of the Mahalanobis distance as a tool for detecting out-of-distribution (OOD) states during inference. Our analysis reveals that while the Mahalanobis distance can successfully identify undesirable states, such as those that lead to the robot becoming stuck, it is not an exact predictor of final task performance. In particular, although the Mahalanobis distance helps measure deviations from the training distribution, it does not always correlate directly with the quantity of rice scooped in our experiments. Nevertheless, these insights underline the potential of the Mahalanobis distance for enhancing policy robustness by flagging critical states during deployment.
@@
Real-Time Operator Takeover Paradigm: We demonstrated how the RTOT framework can improve policy performance by addressing failure scenarios iteratively. The targeted nature of takeover demonstrations provides valuable training data, focusing specifically on challenging states that are often missing from initial demonstrations.
@@
In conclusion, the RTOT paradigm represents a significant step toward more efficient and adaptive robot learning systems. By enabling real-time human intervention and leveraging targeted demonstrations, this approach provides a practical and scalable solution for addressing the inherent challenges of visuomotor policy training, particularly in complex and unpredictable task environments.

### v2

@@ Abstract
We present a Real-Time Operator Takeover (RTOT) paradigm enabling operators to seamlessly take control of a live visuomotor diffusion policy, guiding the system back into desirable states or reinforcing specific demonstrations. We present new insights in using the Mahalonobis distance to automatically identify undesirable states. Once the operator has intervened and redirected the system, the control is seamlessly returned to the policy, which resumes generating actions until further intervention is required. We demonstrate that incorporating the targeted takeover demonstrations significantly improves policy performance compared to training solely with an equivalent number of, but longer, initial demonstrations. We provide an in-depth analysis of using the Mahalanobis distance to detect out-of-distribution states, illustrating its utility for identifying critical failure points during execution. Supporting materials, including videos of initial and takeover demonstrations and all rice scooping experiments, are available on the project website111 https://operator-takeover.github.io/.
@@ I Introduction
While teleoperation systems can boost the performance of IL methods, they do not address the fundamental problem of how to cover sufficient variations in demonstrations/tasks to make the learned policy more robust. Enhancing robustness in IL is an active area of research [15, 16]. Although existing approaches improve the handling of scene variations, they still struggle when encountering states that are not present in the training data. Policies often fail in such scenarios, getting stuck due to the lack of demonstrations for recovering from undesirable states.
In this work, we tackle this challenge with our Real-Time Operator Takeover (RTOT) paradigm. The core idea is that even the most skilled operator may struggle to anticipate and demonstrate recovery scenarios or failure cases the policy might encounter. Our RTOT approach addresses this limitation by first training an initial policy with a smaller set of demonstrations than typically used in the literature. This policy is then deployed in the environment while the operator remains on standby. When the operator observes the system entering or approaching an undesirable state, they seamlessly take control to teleoperate the robot back to a desirable state. This takeover demonstrations are recorded, and a new policy is trained using both the initial and the newly recorded takeover demonstrations. As shown in Fig. 1, this paradigm can be repeated as many times as necessary to achieve the desired performance for any given task. Our contributions are as follows:
@@
    Introduction and validation of the RTOT paradigm on a real-world rice scooping task.
@@ IV Experimental Setup
We evaluate this through a cyclic rice scooping task, as shown in Fig. 1. The objective is to scoop rice from the bowl on the left (positioned in the white rectangle) and deposit it into the bowl on the right (which can be placed anywhere within its respective rectangle). To address the cyclic nature of the task, the initial demonstrations are designed to start and end at approximately the same position after successfully completing one scoop.
To begin, we collect an initial dataset $\mathcal{D}_{I}^{60}$ consisting of 60 expert demonstrations. These demonstrations are performed while varying the locations of the rice bowls to introduce diversity in the training data. From this comprehensive dataset, we extract two smaller subsets: the first 20 demonstrations form $\mathcal{D}_{I}^{20}$, and the first 40 demonstrations form $\mathcal{D}_{I}^{40}$. Using these datasets, we train the corresponding policies: $\pi_{I}^{20}$, $\pi_{I}^{40}$, and $\pi_{I}^{60}$.
Next, we deploy $\pi_{I}^{20}$ in the rice scooping environment as described in Section III. During this deployment, we collect 20 additional takeover demonstrations, denoted as $\mathcal{D}T^{20a}$, while systematically varying the position of the rice bowl. These takeover demonstrations are then combined with the initial dataset $\mathcal{D}I^{20}$ to form an augmented dataset $\mathcal{D}{IT}^{a}$, which is used to train a new policy, $\pi{IT}^{a}$.
@@
Additionally, the combination of initial and takeover demonstrations helps to address failure cases more effectively, since takeover demonstrations specifically focus on challenging scenarios encountered during policy deployment. This iterative process not only improves policy robustness but also reduces the overall training time required to achieve high task performance.
@@ V Experimental Evaluation
We evaluate the performance of five different policies ($\pi_{I}^{20},\pi_{I}^{40},\pi_{IT}^{a},\pi_{I}^{60},\pi_{IT}^{b}$) in the cyclic rice scooping task. The goal of the task is to transfer as much rice as possible from a full bowl to an empty bowl within $45$ seconds, measured in grams. For each policy, we conduct $10$ trials, each with different placement locations for the bowls. Videos of all $50$ experiments are available on the project website.
@@
Results and Analysis. The results reveal a clear trend: the performance of the baseline policies improves with the number of initial demonstrations. For example, $\pi_{I}^{20}$, trained on only $20$ demonstrations, achieved an average of $4.00$ grams of rice scooped across $10$ trials. Increasing the data set to $40$ demonstrations significantly increased the performance, with $\pi_{I}^{40}$ averaging $19.80$ grams. Using all $60$ initial demonstrations, $\pi_{I}^{60}$ achieved a substantial improvement, scooping an average of $41.00$ grams per trial. This highlights the importance of larger datasets to improve baseline performance in visuomotor policy training.
To examine whether real-time operator takeover paradigm enhances task performance, we compare the performance of the baseline $\pi_{I}^{40}$ with that of the takeover-enhanced policy $\pi_{IT}^{a}$. Remarkably, $\pi_{IT}^{a}$ achieved an average of $35.40$ grams scooped, representing a $79\%$ improvement over $\pi_{I}^{40}$. Similarly, the final policy, $\pi_{IT}^{b}$, trained on both initial and two sets of takeover demonstrations, outperformed $\pi_{I}^{60}$ by $20\%$, with an average of $49.20$ grams scooped per trial.
@@ VI To Takeover or Not to Takeover
In the current framework, the decision to initiate a takeover rests solely with the operator. We argue that this is a sensible approach, given that the quality of the initial demonstrations—and more broadly, the overall task performance—is highly influenced by the quality of these demonstrations [5]. However, having a more systematic way of assessing how close the robot is to undesirable states, or detecting when such states occur, could provide significant benefits.
@@
This motivates the need for a metric that evaluates how far observations obtained during inference deviate from the training distribution. This metric could serve as an out-of-distribution (OOD) measure to better inform intervention decisions.
@@ VI-A Mahalanobis Distance as an OOD Measure
We estimate the mean ($\mu$) and covariance matrix ($\Sigma$) of each embedding set using the Minimum Covariance Determinant Estimator implemented in scikit-learn [59]. For each encoded observation $h\in H$, the Mahalanobis distance $d_{M}$ is computed with Algorithm 1.
@@ VI-B Mahalanobis Distance Analysis
However, an interesting anomaly is observed when comparing $\mathcal{D}I^{40}$ to $\mathcal{D}{IT}^{a}$ or $\mathcal{D}I^{60}$ to $\mathcal{D}{IT}^{b}$: the mean Mahalanobis distance does not directly correlate with task performance. For example, $\pi_{IT}^{b}$ outperforms $\pi_{I}^{60}$ in rice scooping but has a similar Mahalanobis distance profile.
Case Study: Trial 3. To explore this further, we analyze a specific trial in Fig. 7. For Trial number three, we compare the Mahalanobis distances of policies $\pi_{I}^{20}$, $\pi_{IT}^{b}$, and $\pi_{I}^{60}$. We observe that $\pi_{I}^{20}$ starts with the highest distance while $\pi_{IT}^{b}$ and $\pi_{I}^{60}$ share similar values (around $40$). The optimal task performance model, $\pi_{IT}^{b}$, exhibits a slight increase in distance as the trial progresses. However, none of the encountered states necessitate takeover, and the distance consistently stays below $100$ throughout the duration of the trial.
The poorly performing policy $\pi_{I}^{20}$ exhibits the highest Mahalanobis distance peaking at around $160$, around the $25s$ mark, when the robot’s spoon moves outside the target bowl and fails to recover. The robot remains stuck in this state until the trial ends, highlighting the relationship between high Mahalanobis distance and undesirable states. In contrast, both $\pi_{I}^{60}$ and $\pi_{IT}^{b}$ maintain relatively lower Mahalanobis distances throughout the trial, with no significant increases. While $\pi_{IT}^{b}$ outperforms $\pi_{I}^{60}$ in task performance, the Mahalanobis distance does not differentiate much between the two policies in this regard. We also observe that the sudden stops, and consequently the camera shaking, significant impact as evidenced by the last peak in $\pi_{IT}^{b}$ (orange) and $\pi_{I}^{60}$ (green) around the 45-second mark, coinciding with the abrupt halt of the robot as the experiment concluded. Additionally, the result indicates that the distance remains constant as long as there are no changes in the image caused by motion.
@@ VII Conclusion
In this work, we introduced a real-time operator takeover (RTOT) paradigm to train visuomotor diffusion policies. This framework enables users to seamlessly take control of the robot from a suboptimal policy, preventing failures during execution. These takeover demonstrations are subsequently integrated into the training process, iteratively refining the policy until the desired performance is achieved.
Our experimental evaluation demonstrates the effectiveness of this approach. Despite the fact that the total time of takeover demonstrations is significantly shorter than that of the initial demonstrations, policies trained on these targeted datasets consistently outperform their counterparts. In a cyclic rice scooping task, policies enhanced with takeover demonstrations achieved superior task performance, highlighting the value of addressing failure cases during policy deployment rather than relying solely on extensive initial datasets.
Beyond showcasing the utility of the RTOT framework, we also explored the use of the Mahalanobis distance as a tool for detecting out-of-distribution (OOD) states during inference. Our analysis reveals that while the Mahalanobis distance can successfully identify undesirable states, such as those that lead to the robot becoming stuck, it is not an exact predictor of final task performance. In particular, although the Mahalanobis distance helps measure deviations from the training distribution, it does not always directly correlate with the quantity of rice scooped in our experiments. Nevertheless, these insights underline the potential of the Mahalanobis distance for enhancing policy robustness by flagging critical states during deployment.
@@
Real-Time Operator Takeover Paradigm: We demonstrated how the RTOT framework can improve policy performance by iteratively addressing failure scenarios. The targeted nature of takeover demonstrations provides valuable training data, focusing specifically on challenging states that are often missing from initial demonstrations.
@@
In conclusion, the RTOT paradigm represents a significant step toward more efficient and adaptive robot learning systems. By enabling real-time human intervention and leveraging targeted demonstrations, this approach provides a practical and scalable solution to address the inherent challenges of visuomotor policy training, particularly in complex and unpredictable task environments.